OpenStudy (pink.1d):

What would a cosine function look like with a given amplitude of 7, period of pi/6, and vertical shift of down 2?

OpenStudy (pink.1d):

please just give me a straighforward answer

OpenStudy (pink.1d):

ill fan and medal:)

OpenStudy (freckles):

\[y=A \cos(b(x-c))+d \\ |A| \text{ is the amp number } \\ \frac{2 \pi}{b} \text{ is the period } \\ c \text{ is the phase shift number } \\ d \text { is the vertical shift } \\ \text{ you are given } |A|=7 \\ \text{ and } \frac{2\pi}{b}=\frac{\pi}{6} \\ \text{ and also what } d \text{ is }\] Now if you want to draw this... Assuming the phase shift is 0.... recall cos(x) is a function between -1 and 1 well you have A cos(x) which puts the function between -A and A A cos(x)+d puts the function between -A+d and A+d So adding A and d and the opposite of A and d give you the max and min of the function recall the special points of y=cos(x) are (0,1) (pi/2,0) (pi,-1) (3pi/2,0) (2pi,1) that is since cos(0)=1 cos(pi/2)=0 cos(pi)=-1 cos(3pi/2)=0 cos(2pi)=1 now since we have the period is a bit difference for our function that will change a bit so say we have y=cos(bx) how does this effect our special numbers we plug in Well the period of y=cos(x) was 2pi and the period for y=cos(bx) is 2pi/b so like we had 4 spaces between our special numbers we will have 4 spaces again 2pi/(4b) so cos(0)=1 cos(0+2pi/(4b))=cos(pi/(2b))=0 cos(pi/(2b)+pi/(2b))=cos(pi/b)=-1 cos(pi/b+pi/(2b))=cos(3pi/(2b))=0 cos(3pi/(2b)+pi/(2b))=cos(2pi/b)=1 so the graph will either be stretched out more or less horizontally so y=cos(2x) looks like  see the period is 2pi/2=pi anyways I hopes this gives you a better idea